OpenStudy (anonymous):

Classify the numbers as rational or irrational. negative 4/5 square root of two 2.5 0.44444444... square root of 16 @Hero @dan815 @nincompoop

OpenStudy (anonymous):

please help me someone

OpenStudy (welshfella):

all rational number can be expressed as a fraction a/b eg 1/2 , 3/4 . 6 ( = 6/1) are all rational

OpenStudy (welshfella):

so how about -4/5 ?

OpenStudy (welshfella):

- thay can be negative fractions as well as posituve

OpenStudy (welshfella):

is -4/5 a fraction?

OpenStudy (anonymous):

so negative 4/5 and 2.5 and the square root of 16 are all rational

OpenStudy (welshfella):

right

OpenStudy (welshfella):

what about 0.4444... this is 0.4 recurring - meaning it 'goes on forever' but can it be written as a fraction?

OpenStudy (anonymous):

no

OpenStudy (welshfella):

it can actually 1/9 = 0.11111111.. so 0.44444444... is 4/9 so its rational

OpenStudy (anonymous):

that makes more sense

OpenStudy (welshfella):

the only one which is irrational is sqrt2

OpenStudy (welshfella):

there are many like that , sqrt3, sqrt6, sqrt8 . sqrt 17 and so on

OpenStudy (anonymous):

ok thanks

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

i have another question

OpenStudy (welshfella):

ok but i only have 5 minutes

OpenStudy (anonymous):

What is the product of 1.7 × 10−13 and 3.5 × 1025? 1.8 × 10^12 2.06 × 10^12 3.5 × 10^12 5.95 × 10^12

OpenStudy (anonymous):

What is the product of 1.7 × 10^−13 and 3.5 × 10^25? 1.8 × 10^12 2.06 × 10^12 3.5 × 10^12 5.95 × 10^12

OpenStudy (welshfella):

multiply 1.7 by 3.5 then add the exponents 10^-13 * 10^25 = 10^(-13+25) = 10^ ?

OpenStudy (welshfella):

gtg

OpenStudy (anonymous):

ok bye